OpenStudy (anonymous):

Find an equation of the tangent line to the graph f at the indicated point. f(x) = xsquared + 1, (2,5)

OpenStudy (ddcamp):

y = x² + 1 dy/dx = 2x x = 2, so dy/dx (the slope) = 4 Plug this into point slope form: y-y₁ = m(x-x₁)

OpenStudy (anonymous):

how do i plug this into point slope form?

OpenStudy (ddcamp):

The point (2,5) and the slope 4 (The slope came from the derivative)

OpenStudy (anonymous):

oh of course! that makes sense...I'm clearly not thinking today. Thanks!

OpenStudy (kenljw):

m = slope of tangent line m =dy/dx= 2x=4 (2,5) 5 = 4 2 + b y=4x-3

OpenStudy (kenljw):

This is a direct consequence that the tangent line must intersect the curve at (2,5)